# Infrequent Resolving Algorithm for Online Linear Programming

### v2

@@ Abstract
Online linear programming (OLP) has gained significant attention from both researchers and practitioners due to its extensive applications, such as online auction, network revenue management and advertising. Existing OLP algorithms fall into two categories: LP-based algorithms and LP-free algorithms. The former one typically guarantees better performance, even offering a constant regret, but requires solving a large number of LPs, which could be computationally expensive. In contrast, LP-free algorithm only requires first-order computations but induces a worse performance, lacking a constant regret bound. In this work, we bridge the gap between these two extremes by proposing an algorithm that achieves a constant regret while solving LPs only $\mathcal{O}(\log\log T)$ times over the time horizon $T$. Moreover, when we are allowed to solve LPs only $M$ times, we propose an algorithm that can guarantee an $\mathcal{O}\left(T^{(1/2+\epsilon)^{M-1}}\right)$ regret. Furthermore, when the arrival probabilities are known at the beginning, our algorithm can guarantee a constant regret by solving LPs $\mathcal{O}(\log\log T)$ times, and an $\mathcal{O}\left(T^{(1/2+\epsilon)^{M}}\right)$ regret by solving LPs only $M$ times. Numerical experiments are conducted to demonstrate the efficiency of the proposed algorithms.
@@ 1 Introduction
Consider a decision maker with $m$ types of resources facing sequentially arriving customers over $T$ discrete time periods. We assume the initial inventory is $T\boldsymbol{\rho}$, where $\boldsymbol{\rho}=(\rho_{1},\dots,\rho_{m})\in\mathbb{R}_{++}^{m}$ is given and denotes the vector of average inventories per period. Each customer arriving at period $t$ is characterized by a reward $r_{t}\in\mathbb{R}_{+}$ and a consumption vector $\boldsymbol{A}_{t}\in\mathbb{R}_{+}^{m}$. The time periods are indexed forward, starting from period $t=1$ and ending at period $t=T$. In each time period $t$, after observing the arriving customer characterized by $(r_{t},\boldsymbol{A}_{t})$, the decision maker must irrevocably decide whether to accept or reject the customer ($x_{t}\in\{0,1\}$) [^1] Our results can be easily extended to the case with multiple actions. without knowing future information. Given full information $\{(r_{t},\boldsymbol{A}_{t})\}_{t=1}^{T}$, we can solve the following (offline) linear program:
@@ 1.1 Literature Review
Unknown Distribution. We start with the stream assuming unknown arrival probabilities. Our work closely relates to studies under the random input assumption, where the coefficients $\left\{(r_{t},\boldsymbol{A}_{t})\right\}_{t=1}^{T}$ are generated i.i.d. from an unknown distribution $\mathcal{P}$. We focus on the asymptotic order of the regret as $T$ increases. Bray (2024) shows that the regret is at least $\mathcal{O}(\log T)$ when the demand distribution has infinite supports. In this paper, we focus on the case with a finite-support demand distribution. We first review the so called LP-based algorithms which require solving many LPs but guarantee good performance. For example, under certain non-degeneracy assumptions, Jasin (2015) proposes an $\mathcal{O}(\log^{2}T)$-regret algorithm that requires solving LPs $\mathcal{O}(\log T)$ times. Subsequently, Chen et al. (2024) consider a similar algorithm that requires solving LPs $T$ times, achieving $\mathcal{O}(1)$ regret under the non-degeneracy assumption and $\mathcal{O}(\sqrt{T}\log T)$ otherwise. Under the non-degeneracy assumption and some distribution assumptions, Li and Ye (2022) provide an $\mathcal{O}(\log T\log\log T)$-regret algorithm from the dual perspective, and the algorithm requires solving LPs $T$ times. Then, under similar assumptions, Ma et al. (2024) provide an $\mathcal{O}(\log T)$-regret algorithm requiring solving LPs $\mathcal{O}(T)$ times. Recently, Xie et al. (2023) remove the non-degeneracy assumption and provide an OLP algorithm with $\mathcal{O}(1)$ regret but still requiring solving LPs $T$ times. Moreover, there are studies on infrequent resolving algorithms under additional assumptions. Given a complete sample path (i.e., $T$ samples), Freund and Banerjee (2019) propose an algorithm with lazy resolving (i.e., $\mathcal{O}(\log\log T)$ expected LP solving times) and prove a constant regret. Given a starting point with an accuracy of $\mathcal{O}(1/T)$, Ma et al. (2024) propose an infrequent resolving algorithm that can achieve $\mathcal{O}(\log T)$ regret under the non-degeneracy assumption by solving LPs $O(\log T)$ times. Both assumptions are stringent when $T$ is large. In this work, without such assumptions, we propose an algorithm that can guarantee a constant regret by solving LPs $\mathcal{O}(\log\log T)$ times.
@@ 2.1 Argmax with Infrequent Resolving (AIR) Policy
We now introduce our policy in Algorithm 1, referred to as the Argmax with Infrequent Resolving (AIR) policy. The time set $\mathcal{T}$ in Algorithm 1 will be specified shortly.
@@
     if $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$ and $u_{j}^{t}\geq d_{j}^{t}-u^{j}_{t}$ then $\triangleright$ Argmax between $u_{j}^{t}$ and $d_{j}^{t}-u_{j}^{t}$
@@
We now explain the intuition of Algorithm 1. The algorithm begins with a predetermined time set $\mathcal{T}$ specifying the time points where resolving is needed. At each time point $t\in\mathcal{T}$, the decision maker updates the empirical estimations $\hat{\boldsymbol{p}}^{t}$ as an approximation to the future demand and solves the corresponding fluid LP based on the current inventory $\boldsymbol{b}^{t}$ and the estimation $\hat{\boldsymbol{p}}^{t}$. The optimal solution $\boldsymbol{y}^{*}$ of the fluid LP provides an approximation to the optimal policy. At each time period $t\notin\mathcal{T}$, the decision maker updates an approximation $\boldsymbol{u}^{t}$ to the optimal solution in a “greedy” fashion as described in the algorithm. We adopt the idea of the “argmax” policy in Arlotto and Gurvich (2019) and Vera and Banerjee (2021) to make accept/reject decisions. At each period $t$, the decision maker observes the arrival type $j$ and accepts the request only if it is feasible to do so (i.e., $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$) and if $u_{j}^{t}\geq d_{j}^{t}-u^{j}_{t}$, where $\boldsymbol{u}_{t}$ approximates the number of customers that should be accepted and $\boldsymbol{d}_{t}$ approximates the future demand. Intuitively, $d_{j}^{t}-u_{t}^{j}$ represents the number of type $j$ requests that should be rejected and the decision maker accepts the request $j$ if more should be accepted than rejected. In the following, we use $\mathcal{A}$ to denote the AIR policy in Algorithm 1.
@@ 3 Known Arrival Probabilities
     if $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$ and $u_{j}^{t}\geq d_{j}^{t}-u^{j}_{t}$ then $\triangleright$ Argmax between $u_{j}^{t}$ and $d_{j}^{t}-u_{j}^{t}$
@@ 4.1 OLP Policy Comparison
AIR AFR ADA SFA* DLD* BUF* AIR AFR ADA SFA* DLD* BUF* AIR
2500 2.5 1.5 7.7 45.6 62.3 48.3 0.062 13.3 13.3 0.002 0.003 0.002 13
5000 2.2 1.2 10.5 57.6 82.6 59.0 0.064 26.4 26.6 0.004 0.005 0.004 13
7500 2.2 1.6 12.0 66.6 96.4 65.7 0.065 39.4 39.3 0.005 0.007 0.005 13
10,000 2.2 1.4 13.2 74.4 109.7 72.5 0.066 54.4 54.2 0.007 0.009 0.007 13
12,500 2.1 1.2 14.3 80.9 118.8 76.0 0.077 66.6 66.5 0.009 0.012 0.009 15
15,000 2.2 1.3 15.3 86.8 128.1 79.7 0.086 73.3 73.2 0.011 0.015 0.011 15
17,500 2.2 1.1 16.6 92.1 136.0 82.9 0.083 84.4 84.3 0.012 0.016 0.012 15
20,000 2.1 1.0 17.4 97.0 141.6 85.9 0.084 103.2 103.1 0.014 0.019 0.013 15
100,000 2.2 - [^4] The case is not studied due to the extremely long computation time. - 192.0 260.1 126.6 0.141 - - 0.080 0.105 0.078 15
200,000 2.1 - - 260.2 330.9 151.6 0.204 - - 0.169 0.219 0.163 15
300,000 2.1 - - 313.8 379.2 166.1 0.231 - - 0.220 0.291 0.213 15

### v3

@@ Abstract
Online linear programming (OLP) has gained significant attention from both researchers and practitioners due to its extensive applications, such as online auction, network revenue management and advertising. Existing OLP algorithms fall into two categories: LP-based algorithms and LP-free algorithms. The former one typically guarantees better performance, even offering a constant regret, but requires solving a large number of LPs, which could be computationally expensive. In contrast, LP-free algorithm only requires first-order computations but induces a worse performance, lacking a constant regret bound. In this work, we study the case where the inputs are drawn from an unknown finite-support distribution, and bridge the gap between these two extremes by proposing an algorithm that achieves a constant regret while solving LPs only $\mathcal{O}(\log\log T)$ times over the time horizon $T$. Moreover, when we are allowed to solve LPs only $M$ times, we propose an algorithm that can guarantee an $\mathcal{O}\left(T^{(1/2+\epsilon)^{M-1}}\right)$ regret. Furthermore, when the arrival probabilities are known at the beginning, our algorithm can guarantee a constant regret by solving LPs $\mathcal{O}(\log\log T)$ times, and an $\mathcal{O}\left(T^{(1/2+\epsilon)^{M}}\right)$ regret by solving LPs only $M$ times. Numerical experiments are conducted to demonstrate the efficiency of the proposed algorithms.
@@ 1 Introduction
Consider a decision maker with $m$ types of resources facing sequentially arriving customers over $T$ discrete time periods. We assume the initial inventory is $T\boldsymbol{\rho}$, where $\boldsymbol{\rho}=(\rho_{1},\dots,\rho_{m})\in\mathbb{R}_{+}^{m}$ is given and denotes the vector of average inventories per period. Each customer arriving at period $t$ is characterized by a reward $r_{t}\in\mathbb{R}_{+}$ and a consumption vector $\boldsymbol{A}_{t}\in\mathbb{R}_{+}^{m}$. The time periods are indexed forward, starting from period $t=1$ and ending at period $t=T$. In each time period $t$, after observing the arriving customer characterized by $(r_{t},\boldsymbol{A}_{t})$, the decision maker must irrevocably decide whether to accept or reject the customer ($x_{t}\in\{0,1\}$) [^1] Our results can be easily extended to the case with multiple actions. without knowing future information. Given full information $\{(r_{t},\boldsymbol{A}_{t})\}_{t=1}^{T}$, we can solve the following (offline) linear program:
@@ 1.1 Literature Review
Unknown Distribution. We start with the stream assuming unknown arrival probabilities. Our work closely relates to studies under the random input assumption, where the coefficients $\left\{(r_{t},\boldsymbol{A}_{t})\right\}_{t=1}^{T}$ are generated i.i.d. from an unknown distribution $\mathcal{P}$. We focus on the asymptotic order of the regret as $T$ increases. Bray (2024) shows that the regret is at least $\Omega(\log T)$ when the demand distribution has infinite supports. In this paper, we focus on the case with a finite-support demand distribution. We first review the so called LP-based algorithms which require solving many LPs but guarantee good performance. For example, under certain non-degeneracy assumptions, Jasin (2015) proposes an $\mathcal{O}(\log^{2}T)$-regret algorithm that requires solving LPs $\mathcal{O}(\log T)$ times. Subsequently, Chen et al. (2024) consider a similar algorithm that requires solving LPs $T$ times, achieving $\mathcal{O}(1)$ regret under the non-degeneracy assumption and $\mathcal{O}(\sqrt{T}\log T)$ otherwise. Under the non-degeneracy assumption and some distribution assumptions, Li and Ye (2022) provide an $\mathcal{O}(\log T\log\log T)$-regret algorithm from the dual perspective, and the algorithm requires solving LPs $T$ times. Then, under similar assumptions, Ma et al. (2024) provide an $\mathcal{O}(\log T)$-regret algorithm requiring solving LPs $\mathcal{O}(T)$ times. Recently, Xie et al. (2023) remove the non-degeneracy assumption and provide an OLP algorithm with $\mathcal{O}(1)$ regret but still requiring solving LPs $T$ times. Moreover, there are studies on infrequent resolving algorithms under additional assumptions. Given a complete sample path (i.e., $T$ samples), Freund and Banerjee (2019) propose an algorithm with lazy resolving (i.e., $\mathcal{O}(\log\log T)$ expected LP solving times) and prove a constant regret. Given a starting point with an accuracy of $\mathcal{O}(1/T)$, Ma et al. (2024) propose an infrequent resolving algorithm that can achieve $\mathcal{O}(\log T)$ regret under the non-degeneracy assumption by solving LPs $O(\log T)$ times. Both assumptions are stringent when $T$ is large. In this work, without such assumptions, we propose an algorithm that can guarantee a constant regret by solving LPs $\mathcal{O}(\log\log T)$ times.
@@ 2.1 Argmax with Infrequent Resolving (AIR) Policy
We now introduce our policy in Algorithm 1, referred to as the Argmax with Infrequent Resolving (AIR) policy. We use $\boldsymbol{e}_{j}$ to denote a vector of zeros except 1 at the $j$-th entry. The time set $\mathcal{T}$ in Algorithm 1 will be specified shortly.
@@
     if $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$ and $u_{j}^{t}\geq d_{j}^{t}-u_{j}^{t}$ then $\triangleright$ Argmax between $u_{j}^{t}$ and $d_{j}^{t}-u_{j}^{t}$
@@
We now explain the intuition of Algorithm 1. The algorithm begins with a predetermined time set $\mathcal{T}$ specifying the time points where resolving is needed. At each time point $t\in\mathcal{T}$, the decision maker updates the empirical estimations $\hat{\boldsymbol{p}}^{t}$ as an approximation to the future demand and solves the corresponding fluid LP based on the current inventory $\boldsymbol{b}^{t}$ and the estimation $\hat{\boldsymbol{p}}^{t}$. The optimal solution $\boldsymbol{y}^{*}$ of the fluid LP provides an approximation to the optimal policy. At each time period $t\notin\mathcal{T}$, the decision maker updates an approximation $\boldsymbol{u}^{t}$ to the optimal solution in a “greedy” fashion as described in the algorithm. We adopt the idea of the “argmax” policy in Arlotto and Gurvich (2019) and Vera and Banerjee (2021) to make accept/reject decisions. At each period $t$, the decision maker observes the arrival type $j$ and accepts the request only if it is feasible to do so (i.e., $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$) and if $u_{j}^{t}\geq d_{j}^{t}-u_{j}^{t}$, where $\boldsymbol{u}_{t}$ approximates the number of customers that should be accepted and $\boldsymbol{d}_{t}$ approximates the future demand. Intuitively, $d_{j}^{t}-u_{j}^{t}$ represents the number of type $j$ requests that should be rejected and the decision maker accepts the request $j$ if more should be accepted than rejected. In the following, we use $\mathcal{A}$ to denote the AIR policy in Algorithm 1.
@@ 3 Known Arrival Probabilities
     if $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$ and $u_{j}^{t}\geq d_{j}^{t}-u_{j}^{t}$ then $\triangleright$ Argmax between $u_{j}^{t}$ and $d_{j}^{t}-u_{j}^{t}$
@@ 4.1 OLP Policy Comparison
AIR AFR ADA SFA* DLD* BUF* AIR AFR ADA SFA* DLD* BUF* AIR AFR ADA
2500 2.5 1.5 7.7 45.6 62.3 48.3 0.062 13.3 13.3 0.002 0.003 0.002 13 2500 2500
5000 2.2 1.2 10.5 57.6 82.6 59.0 0.064 26.4 26.6 0.004 0.005 0.004 13 5000 5000
7500 2.2 1.6 12.0 66.6 96.4 65.7 0.065 39.4 39.3 0.005 0.007 0.005 13 7500 7500
10,000 2.2 1.4 13.2 74.4 109.7 72.5 0.066 54.4 54.2 0.007 0.009 0.007 13 10,000 10,000
12,500 2.1 1.2 14.3 80.9 118.8 76.0 0.077 66.6 66.5 0.009 0.012 0.009 15 12,500 12,500
15,000 2.2 1.3 15.3 86.8 128.1 79.7 0.086 73.3 73.2 0.011 0.015 0.011 15 15,000 15,000
17,500 2.2 1.1 16.6 92.1 136.0 82.9 0.083 84.4 84.3 0.012 0.016 0.012 15 17,500 17,500
20,000 2.1 1.0 17.4 97.0 141.6 85.9 0.084 103.2 103.1 0.014 0.019 0.013 15 20,000 20,000
100,000 2.2 - [^4] The case is not studied due to the extremely long computation time. - 192.0 260.1 126.6 0.141 - - 0.080 0.105 0.078 15 - -
200,000 2.1 - - 260.2 330.9 151.6 0.204 - - 0.169 0.219 0.163 15 - -
300,000 2.1 - - 313.8 379.2 166.1 0.231 - - 0.220 0.291 0.213 15 - -